Will you please show me the steps for getting from a = 500 - (n-1)a/2 to a = 1000/n+1?
Answer by scott8148(6628)

multiplying by 2 ___ 2a = 1000 - (n - 1)a

distributing ___ 2a = 1000 - an + a

subtracting a ___ a = 1000 - an

adding an ___ a + an = 1000

factoring ___ a(n + 1) = 1000

dividing by n+1 ___ a = 1000 / (n + 1)